# Correspondence-Free, Function-Based Sim-to-Real Learning for Deformable Surface Control

### v1

@@ V-B3 Confidence map network $\mathcal{N}_{conf}$
The input of the confidence map network $\mathcal{N}_{conf}$ is a vector comprising the simulated position $\mathbf{p}\in\mathbb{R}^{3}$ of a point on the surface, along with the control points $\mathcal{S}^{c}(\mathbf{a})$ that define the global shape features. This combined input is processed through three fully connected hidden layers, each with 128 neurons and ReLU activation, followed by a fully connected layer with a single neuron and sigmoid activation. The output is a confidence score in the range $[0,1]$.
In our implementation, we first train $\mathcal{N}_{fk}$, and then jointly train $\mathcal{N}_{rbf}$ and $\mathcal{N}_{conf}$ while keeping $\mathcal{N}_{fk}$ unchanged. The RBF-based space warping function $\mathbf{\Phi}$ and the B-spline shape decoder $\mathbf{B}$ are both implemented as differentiable layers, enabling seamless integration into the standard training process of $\mathcal{N}_{rbf}$.
@@ V-C Results of Sim-to-Real Learning by 3D Scanner
In the first study, we apply 10 random unseen actuation parameters to the deformable membrane, the corresponding shapes are predicted by our $\mathcal{N}_{fk}$ and $\mathcal{N}_{s2r}$ networks and compared with the scanned shapes. The results of the comparison are shown in Fig.10, where the geometric errors are visualized on surfaces by colormaps. The statistical error distributions are given as histograms. For the selected two configurations – S1 and S2, the average errors are reduced by 75.5% and 63.3% by using our sim-to-real approach, and the maximal errors are reduced by 41.8% and 58.6% accordingly.
We now study if we can directly learn to predict the deformed shape from the actuation parameters using neural networks. We applied 40 random sets of actuation parameters to deform the membrane and scanned the corresponding physically deformed shapes as the training dataset. The direct learning network consists of two hidden MLP layers with 128 neurons in each layer. The input of the network is actuation parameters (i.e., 9 values for each shape) and the output is 200 points sampled on the deformed surface. Hausdorff distance [37] and chamfer distance [38] are jointly used to reduce both maximum and average errors between two point clouds, with weights of 1.0 and 5.0, respectively. Lipschitz loss[39] is utilized to penalize the Jacobian magnitude of trained neural network. The resultant surfaces of prediction are then generated by RBF-based surface deformation using these 200 predicted sample points. The result comparison between this direct learning method and our method has been given in Fig.10. It can be observed that direct learning leads to results with significantly larger errors.
We further conducted this comparison by using more sample shapes for both the direct learning and our sim-to-real based learning – 500 shapes are captured by using 80% for training and 20% for testing. As the results shown in Fig.11, our methods delivers substantially better performance. When using 40 frames in our sim-to-real based training, the error can already be reduced to a very low level similar to the direct training using 400 frames.
@@ V-C2 Effectiveness of joint learning
We now conduct a study to demonstrate the effectiveness of jointly learning the sim-to-real transfer network and the confidence map network with the help of the differentiable alignment module. The experiments are performed on the soft manipulator with two segments. Due to large deformations, the scanned 3D shapes often contain missing regions (see Fig.14). For comparison, we also generate sim-to-real training results using unchanged correspondences that are pre-determined by a state-of-the-art non-rigid registration method [40]. This actually treats the sample points as ‘virtual’ markers by using $L^{2}$-norm distances. As shown in Fig.14(a), using static correspondences can lead the training to become stuck in poor local minima. Furthermore, we also conduct a test by using a fixed confidence weight $w_{c}(\mathbf{p})=1.0$ but using Chamfer distance. The result is as shown in Fig.14(b), where the surface of transferred model shows unwanted distortion caused by the points falling in the missed regions. In contrast, our joint learning method consistently produces results with significantly smaller geometric errors. Examples of our method applied to different shapes are shown in Fig. 14(b)–(d).
@@ V-C3 Number of samples and kernels
When performing sim-to-real learning with dense point cloud input, we typically downsample the scanned 3D point clouds to a fixed number of points to define $\mathcal{S}^{p}$. To determine an appropriate sampling density, we created datasets with varying numbers of points and trained the sim-to-real network $\mathcal{N}_{s2r}$ on each dataset. After training, these networks were evaluated on unseen cases to assess shape prediction accuracy. As shown in Fig.12(a), the approximation error decreases as the number of sample points increases, and begins to plateau around 1,000 points. Furthermore, when comparing networks trained on sparse but accurately tracked real markers with known correspondences, those trained with correspondence information consistently outperform those trained without it (see Fig.12(b)). This is likely because motion capture datasets provide precise correspondences, whereas 3D scan datasets are subject to alignment noise and inaccuracies.
The number of Gaussian kernels, denoted as $N$ in Eq.(1), is also a factor influencing the accuracy and efficiency of the sim-to-real method. With too few kernels, the space warping function lacks the capacity to capture complex variations, leading to notable errors. Conversely, an excessive number of kernels increases computational demands, heavily impacting the IK computation as it must repeatedly apply the RBF-based space warping. To determine the number of kernels, we maintain 1000 sample points and train different sim-to-real networks on the same dataset with 40 shapes, varying only the number of kernels. Shape approximation errors were then evaluated on four unseen cases, with results shown in Fig.13. As a result, more kernels generally lead to higher prediction accuracy but at the cost of longer computation times. To balance between the accuracy and the efficiency, we usually select 100 kernels for our implementation and tests.
@@ V-C4 Ablation study – geometric regularization
with $\mathbf{c}^{p}(\cdot)$ giving the closest point of $\mathbf{p}_{j}$ on the scanned point cloud $\mathcal{S}^{p}$. The distribution of the distance errors and the normal difference errors are plotted as the histograms shown in Fig.15(b). Note that the histograms of normal errors are capped at 0.06, according to an angular difference of less than $20^{\circ}$ between the two normals. This allows to focus on a range of interest for better comparison.
The experimental tests shown in Fig.15 demonstrate that the geometric regularization loss, derived from the compatibility condition of the RBF warping function (Eq.(7)), is essential for achieving better shape accuracy on the corrected models. These tests also confirm the generality of our network when working on different types of numerical simulators.
@@ V-E2 Motion Planning of Soft Manipulator
The accurate shape prediction in kinematic computation is important for the motion planning of a soft robot. We conducted the physical experiment on a soft manipulator with 6 DoFs as shown in Fig.20. Specifically, a sampling-based motion planning algorithm [43] is employed to compute a sequence of samples as actuation parameters that progressively deform the soft manipulator to touch the target ball while presenting an obstacle stick around the manipulator – see Fig.20(a) for the illustration. Without the sim-to-real correction, the shapes predicted by the kinematic computation are significantly different from the reality. This leads to the risk of collision on a planned collision-free trajectory (see Fig.20(c) for the planned collision-free swept volume and the scanned physical execution colliding with the obstacle stick). The problem can be well solved after incorporating our sim-to-real method into the kinematic computation – see the results in Fig. 20(d).
@@ V-E3 Shape Approximation on Deformable Mannequin
We have tested the performance of our NN-based fast IK solver on a variety of individual shapes presented in the CAESAR dataset. Figure 21 shows the progressive results when applying our IK solver to realize the target shape on the physical setup of pneumatic driven deformable mannequin. It can be observed that our method converges very fast – i.e., around 15 iterations where each iteration takes less than 0.5 second. The distribution of shape approximation errors are generated by scanning the soft mannequin that has been deformed using the resultant actuation determined by our IK solver. Results on three different models are given in Fig.22. The effectiveness of our fast IK solver has been tested on other models randomly selected from the CAESAR dataset as shown in Appendix C.
@@ VI Conclusion and Discussion
In this paper, we addressed the challenge of bridging the gap between simulated and physically deformed free-form surfaces. We proposed a novel sim-to-real framework that jointly learns a deformation function space and a confidence map to transfer simulated geometries to their real-world counterparts using diverse and potentially incomplete input data. Unlike prior methods, our approach does not rely on pre-established correspondences and is robust to partial observations. Integrated into a neural network-based computational pipeline, the method effectively solves the inverse kinematics problem across various pneumatically actuated deformable robots—including a deformable membrane, two soft manipulators, and a deformable mannequin. Our approach advances the state of the art in shape control of deformable free-form surfaces, paving the way for more accurate and practical deployment in soft robotic applications.
@@
Although the experimental tests conducted on four soft robots is promising, our method shares the limitation of all learning-based approaches – the results heavily rely on the quality of a training dataset. Moreover, our method cannot perform well on surfaces with sharp features, which is due to the fact that RBFs are inherently smooth and continuous. Therefore, we only apply it to soft robots with freeform deformable surfaces. In our future work, we will use the soft deformable membrane and mannequin as molds to fabricate products with customized shape. Specifically, as demonstrated in the proof-of-concept experiments in collaboration with the Dutch startup NEFFA [44], our soft robotic mannequin is first deformed to the body shape of a customer and then serves as a reusable 3D mold for making a garment using biodegradable materials (see Fig.23 for the fabrication process). Reliable shape control of deformable soft robots under different temperatures needs to be investigated in our future research. Furthermore, we plan to test the soft manipulators in pick-and-place tasks.
@@ Acknowledgement
The project is partially supported by the chair professorship fund of the University of Manchester and the research fund of UK Engineering and Physical Sciences Research Council (EPSRC) (Ref.#: EP/W024985/1). The sustainable fabrication of garments on a soft deformable mannequin is conducted with the help of Aniela Hoitink and Iman Hadzhivalcheva at NEFFA by using their material and fabrication process.
Appendix A Appendix Title
A-A Algorithm of Fast IK Solver
Here we explain the details about how to determine the solution of IK by an iterative algorithm using the computational pipeline of neural network-based forward kinematics, which has integrated the sim-to-real module.
Starting from the initial state, in each iteration, we alternately update the transformation $\mathbf{T}_{i}$ (applied on the target shape $\mathcal{S}^{t}$ by ICP) and the actuation parameters $\mathbf{a}_{i}$. The actuation update process employs a gradient-based method with line search. These approaches gradually drive the deformable free-form surface to approximate the target shape. The pseudo-code outlining this algorithm can be found in Alg. 1, and a corresponding diagram has been provided in Fig. 7.
1
Output: The optimized actuation $\mathbf{a}_{opt}$.
2
3Set $i=0$ and $r_{0}=1.0$;
6while $i<i_{max}$ and $r_{i}>\lambda$ do
 /* Update transformation matrix */
 /* Gradient-based iteration with fixed transformation matrix */
8   Evaluate the objective function $D_{i}=D(\mathbf{a}_{i})$;
9  Compute the gradient of $D(\mathbf{a}_{i})$ as $\nabla D$ by differentiating NN-based Forward Kinematics Pipeline as shown in Eq. (IV-B);
10  Compute the optimal step size $h$ by line search [45];
11  Set $\mathbf{a}_{i+1}=\mathbf{a}_{i}-h\nabla D$ and $i=i+1$;
12  Evaluate the objective function $D_{i}=D(\mathbf{a}_{i})$;
[…]
14 end while
15return $\mathbf{a}_{opt}=\mathbf{a}_{i}$;
ALGORITHM 1 Fast IK Solver
A-B Differentiation of NN-based Forward Kinematics
@@
Note that the last term of Eq. (19) is similar to Eq. (A-B), and we only need to change the sign as
@@
The terms $\frac{\partial\mathbf{Q}}{\partial\mathcal{N}_{fk}}$ and $\frac{\partial\mathbf{B}}{\partial\mathcal{N}_{fk}}$ are the gradients of the point(s) on the simulation surface with respect to control points $\mathcal{S}^{c}$ (stored as a flattened column vector). According to the equation of a B-spline surface [46]:
$\mathbf{B}(u,v)=\sum_{i=1}^{M_{u}}\sum_{j=1}^{M_{v}}N_{ik}(u)N_{jl}(v)\mathcal{S}^{c}_{ij},$ (24)
the sample point $\mathbf{B}(u,v)$ on the B-spline surface is the sum of the linear combination of the B-spline basis function ($N_{ik}(u)$ and $N_{jl}(v)$) and control point $\mathcal{S}^{c}_{ij}$. $k$ and $l$ denote the order of B-spline basis function in $u$ and $v$ direction (cubic B-spline is chosen in the paper). $M_{u}=30$ and $M_{v}=30$ are the control point numbers along $u$ and $v$ direction (i.e., the total number is $900$). The gradient can then be calculated through:
$\frac{\partial\mathbf{B}}{\partial\mathcal{S}^{c}_{ij}}=N_{ik}(u)N_{jl}(v)\mathbf{I}$ (25)
@@
A-C Additional IK Tests with More Target Shapes
@@
Figure 24: To verify the effectiveness of our fast IK solver, we have selected 10 target models exhibiting a wide range of body dimensions. These models encompass those previously discussed in our paper (denoted as M1M4M1-M4) as well as six additional models denoted as S1S6S1-S6. It can be found that very small errors of shape approximation are generated between the physical results and the target shapes.

### v2

@@ V-B3 Confidence map network $\mathcal{N}_{conf}$
The input of the confidence map network $\mathcal{N}_{conf}$ is a vector comprising the simulated position $\mathbf{p}\in\mathbb{R}^{3}$ of a point on the surface, along with the control points $\mathcal{S}^{c}$ that define the global shape features. This combined input is processed through three fully connected hidden layers, each with 128 neurons and ReLU activation, followed by a fully connected layer with a single neuron and sigmoid activation. The output is a score in the range of $[0,1]$.
In our implementation, we first train $\mathcal{N}_{fk}$, and then jointly train $\mathcal{N}_{rbf}$ and $\mathcal{N}_{conf}$ while keeping $\mathcal{N}_{fk}$ unchanged. The RBF-based space warping function $\mathbf{\Phi}$ and the B-spline shape decoder $\mathbf{B}$ are both implemented as differentiable layers, enabling seamless integration into the training process of $\mathcal{N}_{rbf}$ by backpropagation.
@@ V-C Results of Sim-to-Real Learning by 3D Scanner
In the first study, we apply 10 random unseen actuation parameters to the deformable membrane, where the corresponding shapes are predicted by our $\mathcal{N}_{fk}$ and $\mathcal{N}_{s2r}$ networks and compared with the scanned shapes. The results of the comparison are shown in Fig.10, where the geometric errors are visualized on surfaces by colormaps. The statistical error distributions are given as histograms. For the selected two configurations – S1 and S2, the average errors are reduced by 75.5% and 63.3% when using our sim-to-real approach, and the maximal errors are reduced by 41.8% and 58.6% accordingly.
An experiment is taken now to study if we can directly learn to predict the deformed shape from the actuation parameters using neural networks. We applied 40 random sets of actuation parameters to deform the membrane and then scanned the corresponding physically deformed shapes as the training dataset. The direct learning network consists of two hidden MLP layers with 128 neurons in each layer. The input of the network is actuation parameters (i.e., 9 values for each shape) and the output is 200 points sampled on the deformed surface. To reduce the geometry discrepancy between the predicted surfaces $\mathcal{S}_{j}^{*}$ and the scanned surfaces $\mathcal{S}_{j}^{p}$, the Hausdorff distance [39] as
$D_{\text{Haus}}(\mathcal{S}_{j}^{*},\mathcal{S}_{j}^{p})$ $=\max\Big(\max_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\min_{\mathbf{x}\in\mathcal{S}_{j}^{p}}\|\mathbf{p}^{*}-\mathbf{x}\|_{2},$
$\quad\;\;\max_{\mathbf{x}\in\mathcal{S}_{j}^{p}}\min_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\|\mathbf{p}^{*}-\mathbf{x}\|_{2}\Big),$ (18)
and the chamfer distance [40] as
$D_{\text{Cham}}(\mathcal{S}_{j}^{*},\mathcal{S}_{j}^{p})$ $=\frac{1}{|\mathcal{S}_{j}^{*}|}\sum_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\min_{\mathbf{x}\in\mathcal{S}_{j}^{p}}\|\mathbf{p}^{*}-\mathbf{x}\|_{2}^{2}$
$\quad+\frac{1}{|\mathcal{S}_{j}^{p}|}\sum_{\mathbf{x}\in\mathcal{S}_{j}^{p}}\min_{\mathbf{p}^{*}\in\mathcal{S}_{j}^{*}}\|\mathbf{p}^{*}-\mathbf{x}\|_{2}^{2}$ (19)
are jointly used to penalize the maximum and the average deviations between two point clouds by a loss as
$\color[rgb]{0,0,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0.0}\mathcal{L}_{direct}=D_{\text{Haus}}+5.0D_{\text{Cham}}.$}} (20)
The resultant surfaces of prediction are then generated by RBF-based surface deformation using these 200 predicted sample points. The result comparison between this direct learning method and our method has been given in Fig.10. It can be observed that direct learning leads to results with significantly larger errors.
We further conducted this comparison by using more sample shapes for both the direct learning and our sim-to-real based learning – 500 shapes are captured by using 80% for training and 20% for testing. As shown in Fig.11, our method also delivers substantially better performance. When using 40 frames in our sim-to-real based training, the error can already be reduced to a very low level similar to the direct training using 400 frames.
@@ V-C2 Effectiveness of joint learning
We now conduct a study to demonstrate the effectiveness of jointly learning the sim-to-real transfer network and the confidence map network with the help of the differentiable alignment module. The experiments are performed on the soft manipulator with two segments. Due to large deformations, the scanned 3D shapes often contain missing regions (see Fig.12). For comparison, we also generate sim-to-real training results using unchanged correspondences that are pre-determined by a state-of-the-art non-rigid registration method [41]. This actually treats the sample points as ‘virtual’ markers by using $L^{2}$-norm distances. As shown in Fig.12(a.1), using static correspondences can lead the training to become stuck in poor local minima. Furthermore, we also conduct a test by using a fixed confidence weight $w_{c}(\mathbf{p})=1.0$ but using Chamfer distance. The result is as shown in Fig.12(a.2), where the surface of transferred model shows unwanted distortion caused by the points falling in the missed regions. In contrast, our joint learning method consistently produces results with significantly smaller geometric errors. Examples of our method applied to different shapes are shown in Fig. 12(b)–(d).
@@ V-C3 Number of samples and kernels
When performing sim-to-real learning with dense point cloud input, we usually downsample the scanned 3D point clouds to a fixed number of points to define $\mathcal{S}^{p}$. To determine an appropriate sampling density, we created datasets with varying numbers of points and trained the sim-to-real network $\mathcal{N}_{s2r}$ on each dataset. After training, these networks were evaluated on unseen cases to assess shape prediction accuracy. As shown in Fig.13(a), the approximation error decreases as the number of sample points increases, and begins to plateau around 1,000 points. Furthermore, when comparing networks trained on point clouds without correspondence, those incorporating MoCap-provided correspondences consistently achieve higher accuracy under comparable sampling densities (see Fig.13(b)). This is likely because motion capture datasets provide precise correspondences, whereas 3D scan datasets are subject to alignment noise and inaccuracies.
The number of Gaussian kernels, denoted as $N$ in Eq.(2), is also a factor influencing the accuracy and efficiency of the sim-to-real method. With too few kernels, the space warping function lacks the capacity to capture complex variations, leading to notable errors. Conversely, an excessive number of kernels increases computational demands, heavily impacting the IK computation as it must repeatedly apply the RBF-based space warping. To determine the number of kernels, we maintain 1000 sample points and train different sim-to-real networks on the same dataset with 40 shapes, varying only the number of kernels. Shape approximation errors were then evaluated on four unseen cases, with results shown in Fig.14. As a result, more kernels generally lead to higher prediction accuracy but come with the cost of longer computation times. To balance between the accuracy and the efficiency, we usually select 100 kernels for our implementation and tests.
@@ V-C4 Ablation study – geometric regularization
with $\mathbf{c}^{p}(\cdot)$ giving the closest point of $\mathbf{p}_{j}$ on the scanned point cloud $\mathcal{S}^{p}$. The distribution of the distance errors and the normal difference errors are plotted as the histograms shown in Fig.15(b). Note that the histograms of normal errors are capped at 0.06 – i.e., an angular difference of less than $20^{\circ}$ between the two normals. This allows to focus on a range of interest for better comparison.
The experimental tests shown in Fig.15 demonstrate that the geometric regularization loss, derived from the compatibility condition of the RBF warping function (i.e., Eq.(10)), is essential for achieving better shape accuracy on the corrected models. These tests also confirm the generality of our network when working on different types of numerical simulators.
@@ V-E2 Motion Planning of Soft Manipulator
The accurate shape prediction in kinematic computation is important for the motion planning of a soft robot. We conducted the physical experiment on a soft manipulator with 6 DoFs as shown in Fig.20. Specifically, a sampling-based motion planning algorithm [44] is employed to compute a sequence of samples as actuation parameters that progressively deform the soft manipulator to touch the target ball while presenting an obstacle stick around the manipulator – see Fig.20(a) for the experimental setup. Without the sim-to-real correction, the shapes predicted by the kinematic computation are significantly different from the reality. This leads to the risk of collision on a planned collision-free trajectory (see Fig.20(c) for the planned collision-free swept volume and the scanned physical execution colliding with the obstacle stick). The problem can be well solved after incorporating our sim-to-real method into the kinematic computation – see the results in Fig. 20(d).
@@ V-E3 Shape Approximation on Deformable Mannequin
We have tested the performance of our NN-based fast IK solver on a variety of individual shapes presented in the CAESAR dataset [45]. Fig.21 shows the progressive results when applying our IK solver to realize the target shape on the physical setup of pneumatic driven deformable mannequin. It can be observed that our method converges very fast – i.e., around 15 iterations where each iteration takes less than 0.5 second. The distribution of shape approximation errors are generated by scanning the soft mannequin that has been deformed using the resultant actuation determined by our IK solver. Results on three different models are given in Fig.22. Our fast IK solver has also been tested on other models randomly selected from the CAESAR dataset as shown in Appendix A-B.
@@ VI Conclusion and Discussion
In this paper, we addressed the challenge of bridging the gap between simulated and physically deformed free-form surfaces. We proposed a novel sim-to-real framework that jointly learns a deformation function space and a confidence map to transfer simulated geometries to their real-world counterparts using diverse and potentially incomplete input data. Unlike prior methods, our approach does not rely on pre-established correspondences and is robust to partial observations. Integrated into a neural network-based computational pipeline, the method effectively solves the inverse kinematics problem across various pneumatically actuated deformable robots, including a deformable membrane, two soft manipulators, and a deformable mannequin. Our approach advances the state of the art in shape control of deformable free-form surfaces, paving the way for more accurate and practical deployment in soft robotic applications.
@@
Although the experimental tests conducted on four soft robots is promising, our method shares the limitation of all learning-based approaches – the results heavily rely on the quality of training datasets. Moreover, our method cannot perform well on surfaces with sharp features, which is due to the fact that RBFs are inherently smooth and continuous. Therefore, we only apply it to soft robots with freeform deformable surfaces. In our future work, we will use the soft deformable membrane and mannequin as molds to fabricate products with customized shape. Specifically, as demonstrated in the proof-of-concept experiments in collaboration with the Dutch startup NEFFA [46], our soft robotic mannequin is first deformed to the body shape of a customer and then serves as a reusable 3D mold for making a garment using biodegradable materials (see Fig.23 for the fabrication process). Reliable shape control of deformable soft robots under different temperatures needs to be investigated in our future research. Furthermore, we plan to test the soft manipulators in pick-and-place tasks.
@@ Acknowledgement
The project is partially supported by the chair professorship fund of the University of Manchester and the research fund of UK Engineering and Physical Sciences Research Council (EPSRC) (Ref.#: EP/W024985/1). The sustainable fabrication of garments shown in Fig.23 is conducted with the help of Aniela Hoitink and Iman Hadzhivalcheva at NEFFA by using their material and fabrication process.
Appendix A Differentiation of NN-based Forward Kinematics
A-A Differentiation of NN-based Forward Kinematics
@@
Note that the last term of Eq. (26) is similar to Eq. (A-A), and we only need to change the sign as
@@
The terms $\frac{\partial\mathbf{Q}}{\partial\mathcal{N}_{fk}}$ and $\frac{\partial\mathbf{B}}{\partial\mathcal{N}_{fk}}$ are the gradients of the point(s) on the simulation surface with respect to control points $\mathcal{S}^{c}$ (stored as a flattened column vector). According to Eq.(1), the gradient can then be calculated through:
$\frac{\partial\mathbf{B}}{\partial\mathcal{S}^{c}_{ij}}=N_{ik}(u)N_{jl}(v)\mathbf{I}.$ (31)
@@
A-B Additional IK Tests with More Target Shapes